OpenStudy (anonymous):

3^2+b^2=4^2 I ended up with b^2=7 b=sq rt 7

OpenStudy (amistre64):

+ or - sqrt()7

OpenStudy (amistre64):

(-x)^2 = x^2 (x)^2 = x^2 + or - for those

OpenStudy (anonymous):

yes u r absolutely correct keep it up : transpose 2^2 to right hand side b^2 = 4^2-3^22 \[b^2=4^2-3^2\] \[b^2=16-9\] or \[b^2=(4+3)(4-3)\] \[b^2=7\] \[b=sqrt7\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

a medal would be gr8 kylie